My mothers age is three times my age. The sum of our ages is 40. How old am I? How old is my mom? Let x be your age.
Answer by ankor@dixie-net.com(22740)

My mothers age is three times my age.
The sum of our ages is 40.
How old am I? How old is my mom?
Let x be your age.
;
"My mothers age is three times my age."
m = 3x
:
"The sum of our ages is 40."
m + x = 40
Replace m with 3x
3x + x = 40
4x = 40
x = 40/4
x = 10
:
I'll let you find Mom's age